measure of angle T is 20 more than four times its supplement
Answer by mananth(16949)

let supplement angle be x
T= 4x+20
..
x+4x+20=180
5x+20=180
add -20
5x+20-20=180-20
5x=160
divide by 5
x=160/5
x=32 deg
T= 4*x+20
T=4*32+20
T=128+20
T=148 deg